# OWT: A Foundational Organ-Wise Tokenization Framework for Medical Imaging

### v1

@@ 5 Conclusion
OWT/${}_{\text{4slices}}$ 96.81 94.84 95.59 78.23 91.37 +2.03
AutoPet Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus [63] 96.41 92.56 91.31 76.87 89.29
AttU_Net [42] 96.60 92.86 91.22 78.24 89.73
TransUnet [9] 96.34 93.59 90.65 78.23 89.70
SwinUnet [5] 96.33 92.34 91.11 72.46 88.06
CMUNet [51] 96.56 93.11 90.61 77.65 89.48
@@
OWT/${}_{\text{4slices}}$ 96.65 93.15 93.96 81.87 91.41 +1.68
Table 7: Performance of semantic segmentation on CT datasets.
Delay Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus [63] 93.53 91.85 85.20 65.85 84.11
AttU_Net [42] 94.25 91.08 86.85 65.98 84.54
TransUnet [9] 93.93 90.83 89.07 66.03 84.97
SwinUnet [5] 92.78 90.61 86.35 62.55 83.07
CMUNet [51] 94.02 91.88 87.74 65.64 84.82
[…]
OWT/${}_{\text{4slices}}$ 94.72 92.61 92.23 71.33 87.72 +1.87
PreArtery Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus [63] 94.05 92.25 87.26 64.14 84.42
AttU_Net [42] 94.33 92.32 89.07 64.45 85.04
TransUnet [9] 94.33 90.69 86.59 65.56 84.29
SwinUnet [5] 92.91 88.91 86.33 61.48 82.41
CMUNet [51] 94.26 92.58 89.21 66.18 85.56
[…]
OWT/${}_{\text{4slices}}$ 94.74 92.18 92.14 70.17 87.31 +1.64
Table 8: Performance of semantic segmentation on MRI datasets.
6 Implementation Details
6.1 Image Preprocessing
For all 3D medical volumes, we standardize the spatial resolution to $2\text{mm}\times 2\text{mm}\times 2\text{mm}$ and apply a center crop of $224\times 224$. For CT datasets, we clip intensity values to the $[-1000,1000]$ Hounsfield Unit (HU) range and normalize them to $[0,1]$. For MRI datasets, we clip intensities to the $[0,99]$ percentile range and normalize them to $[0,1]$. We define five token groups to extract semantic features based on common annotations: background, liver, kidneys, spleen, and pancreas. Each medical image is further standardized in thickness by center-padding or cropping 112 slices based on the anatomical centers of the four organs. For the Abdomen1K dataset [40], the number of cases in the training and test sets is 899:101. For AutoPet [21], the split is 700:200, while for RAOS delay and pre-artery sequences [39], both have a training-to-test ratio of 330:83. All experiments, including OWT, MAE, and segmentation networks, are based on the same preprocessed dataset.
For the 2D face images in CelebAMaskHQ dataset [34], we resize them to $224\times 224$, and define six semantic entities based on annotations (background, face&ear, hair&hat, mouth, nose, and eyes&glasses), as shown in Fig. 8. The CelebAMaskHQ dataset contains 30,000 facial image cases, with 24,000 images allocated for training and 6,000 for testing.
6.2 Architecture Details
6.2.1 Default Architecture of MAE and OWT
In OWT, we adopt the same encoder and decoder architecture as MAE, with the only modification being the replacement of the self-attention blocks [53] with linear self-attention blocks [31]. By default, we use the ViT-base architecture, where the MAE encoder consists of 12 attention blocks, the decoder has 8, and the token embedding dimension is set to 768.
For OWT, we configure the encoder and Token Group Encoder with 6 attention blocks each, while keeping the decoder at 8 blocks. This ensures that the overall model length remains consistent with MAE. Additionally, since each SDTG contains a default of 20 tokens, the computational load of OWT is lower than that of MAE when using fewer than 10 SDTGs. To prevent information leakage during TGR training, we remove the class token from the OWT architecture. For OWT processing 2D images, we use a $16\times 16$ kernel for patch embedding. For 3D input volumes, we apply a $1\times 16\times 16$ 3D kernel for patch embedding.
6.2.2 OWT w/ Text Encoding
For the OWT experiment incorporating text encoding, we design a text prompt inspired by [35]. Specifically, we define the template as "a computerized tomography [category]", where [category] corresponds to "background without the Liver, Kidneys, Spleen, and Pancreas", "of a Liver", "of Kidneys", "of a Spleen", and "of a Pancreas", aligning with the number of token groups. These prompts are fed into a pretrained CLIP text encoder [45] to obtain the corresponding text encoding, resulting in a representation of size $5\times 512$.
To match the token groups, we replicate each text encoding according to the number of tokens per token group (20), leading to a final text encoding dimension of $100\times 512$. Within the OWT framework, a linear transformation layer maps the text encoding from $512$ to $768$, producing a transformed text embedding of size $100\times 768$. This transformed text encoding is then element-wise added to the organ-wise representation $X_{G}$ extracted by the Organ Collector. The combined representation proceeds through the random selection procedure in TGR, while all other components remain consistent with the default OWT settings.
6.3 Experimental Setup
[…]
@@ 6.3 Experimental Setup
7 Detailed Semantic Segmentation Results
Detailed Results. Tables 7 and 8 provide a more detailed results of those presented in Table 3, showing the Dice scores for each of the four organs across different datasets. From Tables 7 and 8, we observe that compared to dedicated segmentation networks, OWT/${}_{\text{4slices}}$ generally achieves superior performance across all organs. Notably, for smaller organs such as the pancreas, OWT/${}_{\text{4slices}}$ demonstrates a significant improvement, with a Dice score advantage exceeding 5.3 in the best case.
@@ 7 Detailed Semantic Segmentation Results
Method AutoPet CelebAMaskHQ
L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
MAE/${}_{\text{LPIPS}}$ 2.77e-04 0.0255 2.78e-03 0.1157
OWT 1.49e-04 0.0232 2.50e-03 0.1101
OWT/${}_{\text{VQGAN}}$ 6.89e-05 ↓53.8% 0.0073 ↓68.5% 1.80e-03 ↓43.8% 0.0540 ↓51.0%
8 Ablation Studies
For the ablation study, we explore the following key aspects: the flexibility of the encoder-decoder architecture, the number of tokens per SDTG, and the impact of varying the amount of input data in OWT.
Backbone Alteration. We investigate the effect of replacing the default ViT-based encoder-decoder architecture in OWT (adopted from MAE) with the encoder and decoder from VQGAN. Specifically, we configure the VQGAN bottleneck dimensions to $14\times 14$ for a $224\times 224$ input image, ensuring that the number of latent tokens (196) remains consistent with the ViT-based setting. As shown in Table 9, despite the increased computational cost of VQGAN’s hybrid CNN and self-attention structure, it achieves significant performance gains over the transformer-based encoder-decoder. On two (medical and facial) datasets, it improves both L2 and LPIPS evaluation metrics by 43.8%–68.5%, demonstrating the potential for further enhancements in OWT and highlighting its architectural flexibility.
Number of Tokens per SDTG. The number of tokens within each SDTG determines its capacity to store organ-specific information. To evaluate this, we experiment with SDTG token counts of 5, 10, 20, and 40, assessing their impact on both generation and segmentation tasks. For the generation task, we measure full-image reconstruction using all SDTGs and single-organ reconstruction using only one SDTG, evaluated by L2, LPIPS, and SSIM-3D metrics. For segmentation, we assess the Dice score of single-organ masks.
Data Percentage. Fig. 12 illustrates the performance of the OWT framework on the AutoPet dataset when trained with 20%, 60%, and 100% of the data. We observe a significant improvement in performance when increasing the dataset size from 20% to 60%, whereas the gain from 60% to 100% is relatively marginal. This suggests that OWT efficiently learns meaningful representations with a moderate amount of data, indicating its strong data efficiency and the potential for effective generalization even with limited training samples.

### v2

@@ B.2 Default Architecture of OWT and MAE
OWT 1.49e-04 0.0232 2.50e-03 0.1101
@@
B.3 Experimental Setup
We follow the same experimental setup and hyperparameter settings as MAE. Additionally, each MAE- and OWT-based experiment is trained for 1200 epochs, with a 60-epoch warm-up phase. The training follows the learning rate protocol $\text{lr}=\text{base\_lr}\times{\text{effective\_batch\_size}}/{256}$ where the base learning rate is set to $\text{base\_lr}=10^{-4}$.
@@ B.3 Experimental Setup
Appendix C Detailed Settings of Multi-Modality Incorporation
For the OWT experiment incorporating text encoding, we design a text prompt inspired by [liu2023clip]. Specifically, we define the template as "a computerized tomography [category]", where [category] corresponds to "background without the Liver, Kidneys, Spleen, and Pancreas", "of a Liver", "of Kidneys", "of a Spleen", and "of a Pancreas", aligning with the number of token groups. These 5 prompts are fed into a CT-specific pretrained CLIP text encoder, CT-CLIP [hamamci2024developing] to obtain the corresponding CT-consistent text encoding, resulting in a representation of size $5\times 512\times 768$. Then, the mean is taken along the middle dimension, yielding a representation of size $5\times 768$.
To match the dimension token groups, we replicate each text encoding according to the number of tokens per token group (20), leading to a final text encoding dimension of $100\times 768$. Within the OWT framework, a linear transformation layer maps the text encoding from $768$ to $768$, resulting in a transformed text embedding of the same size $100\times 768$. This transformed text encoding is then element-wise added to the organ-wise representation $X_{G}$ extracted by the Organ Collector. The combined representation proceeds through the random selection procedure in TGR, while all other components remain consistent with the default OWT settings.
Appendix D Ablation Studies
For the ablation study, we explore the following key aspects: the flexibility of the encoder-decoder architecture, the number of tokens per SDTG, and the impact of varying the amount of input data and labels in OWT.
D.1 Backbone Alteration
We investigate the effect of replacing the default ViT-based encoder-decoder architecture in OWT (adopted from MAE) with the encoder and decoder from VQGAN [esser2021taming]. Specifically, we configure the VQGAN bottleneck dimensions to $14\times 14$ for a $224\times 224$ input image, ensuring that the number of latent tokens (196) remains consistent with the ViT-based setting. As shown in Table 11, despite the increased computational cost of VQGAN’s hybrid CNN and self-attention structure, it achieves significant performance gains over the transformer-based encoder-decoder. On two datasets, it improves both L2 and LPIPS evaluation metrics by 53.8%–68.5% in AutoPet and 43.8%–51.0% in CelebAMaskHQ, demonstrating the potential for further enhancements in OWT and highlighting its architectural flexibility.
AutoPet Holistic Rec. 1 Token Group Rec.
OWT L2 $\downarrow$ LPIPS $\downarrow$ L2 $\downarrow$ LPIPS $\downarrow$
Regular 1.49e-04 0.0232 3.32e-04 0.0071
Adaptive 1.40e-04 0.0215 3.19e-04 0.0067
Table 12: Reconstruction comparison of OWT with regular and adaptively distributed SDTGs.
D.2 Number of Tokens per SDTG
The number of tokens within each SDTG determines its capacity to store organ-specific information. To evaluate this, we experiment with SDTG token counts of 5, 10, 20, and 40, assessing their impact on both generation and segmentation tasks. For the generation task, we measure full-image reconstruction using all SDTGs and single-organ reconstruction using only one SDTG, evaluated by L2, LPIPS, and SSIM-3D metrics. For segmentation, we assess the Dice score of single-organ masks.
Abdomen1k Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus 96.85 94.63 91.00 71.05 88.38
AttU_Net 97.09 95.22 93.05 72.00 89.34
TransUnet 96.78 94.65 92.30 71.89 88.91
SwinUnet 96.53 93.68 90.52 65.50 86.56
CMUNet 96.05 94.67 90.84 71.75 88.33
OWT 96.55 94.21 95.12 71.32 89.30
[…]
AutoPet Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus 96.41 92.56 91.31 76.87 89.29
AttU_Net 96.60 92.86 91.22 78.24 89.73
TransUnet 96.34 93.59 90.65 78.23 89.70
SwinUnet 96.33 92.34 91.11 72.46 88.06
CMUNet 96.56 93.11 90.61 77.65 89.48
OWT 96.14 92.33 92.79 76.82 89.52
[…]
Table 13: Performance of semantic segmentation on CT datasets.
Delay Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus 93.53 91.85 85.20 65.85 84.11
AttU_Net 94.25 91.08 86.85 65.98 84.54
TransUnet 93.93 90.83 89.07 66.03 84.97
SwinUnet 92.78 90.61 86.35 62.55 83.07
CMUNet 94.02 91.88 87.74 65.64 84.82
OWT 94.32 91.87 91.55 65.65 85.85
[…]
PreArtery Liver Kidney Spleen Pancreas Avg Dice $\uparrow$
UNetplus 94.05 92.25 87.26 64.14 84.42
AttU_Net 94.33 92.32 89.07 64.45 85.04
TransUnet 94.33 90.69 86.59 65.56 84.29
SwinUnet 92.91 88.91 86.33 61.48 82.41
CMUNet 94.26 92.58 89.21 66.18 85.56
OWT 94.19 91.10 91.82 65.57 85.67
[…]
Table 14: Performance of semantic segmentation on MRI datasets.
D.3 Adaptively Distributed SDTGs
To explore this, we include a proof-of-concept experiment showing that the OWT framework has the potential to support adaptive token group sizing. We allocated token numbers to different organs based on the fourth root of their volumes. For background (containing all information other than defined organs), liver, kidney, spleen, and pancreas, 45, 20, 13, 13, and 9 were allocated, with the total tokens fixed at 100 (identical to the default setting). Preliminary results show that this adaptive allocation better accommodates the varying feature distributions across organs and provides performance improvements over a fixed-size setting (Table 12), demonstrating the scalability of OWT as a novel foundational tokenization framework.
D.4 Data Percentage
Fig. 12 illustrates the performance of the OWT framework on the AutoPet dataset when trained with 20%, 60%, and 100% of the data. We observe a significant improvement in performance when increasing the dataset size from 20% to 60%, whereas the gain from 60% to 100% is relatively marginal. This suggests that OWT efficiently learns meaningful representations with a moderate amount of data, indicating its strong data efficiency and the potential for effective generalization even with limited training samples.
Figure 13: Semantic-based retrieval of CT images.
Appendix E Detailed Semantic Segmentation Results
[…]
@@ Appendix E Detailed Semantic Segmentation Results
Appendix F Additional Organ-Level Retrieval Demonstration
An additional example demonstrates OWT performing organ-level retrieval on the CT modality (Fig. 13). Combined with Fig. 5, this shows the stability and potential of OWT for organ-level retrieval across different modalities. It can not only retrieve patients with similar body shapes at a holistic level, but also retrieve organs in similar conditions and locations for specific organs of interest. This highlights the more flexible and unique clinical value of OWT, which cannot be achieved by existing retrieval methods based on holistic embeddings.